### v4

@@ A.5 Proof in Section 7
$\max_{x\in\mathbb{R}}w^{n}(x)$ $=w^{n}(\overline{x})\leq\frac{\exp\left(-\frac{c}{\lambda_{1}}\right)}{r+\exp\left(-\frac{c}{\lambda_{1}}\right)}(\mathcal{M}^{\lambda_{2}}\psi^{\bm{\lambda},n}(\overline{x})-\mathcal{M}^{\lambda_{2}}\psi^{\bm{\lambda},n-1}(\overline{x}))$ (A.17)
$\leq\left(1-\frac{r}{r+C_{0}}\right)\|\mathcal{M}^{\lambda_{2}}\psi^{\bm{\lambda},n}-\mathcal{M}^{\lambda_{2}}\psi^{\bm{\lambda},n-1}\|_{L^{\infty}(\mathbb{R})}$
$\leq\frac{C_{0}}{r+C_{0}}\|\psi^{\bm{\lambda},n}-\psi^{\bm{\lambda},n-1}\|_{L^{\infty}(\mathbb{R})},$
where the last inequality is due to the non-expansiveness of $\mathcal{M}^{\lambda_{2}}$ by Lemma 3.1. Following a similar line of proof, we assume that $w^{n}$ attains its minimum at $\underline{x}$, then
$0$ $\leq\mathcal{L}w^{n}(\underline{x})$
$=\left(r+\exp\left(-\frac{c}{\lambda_{1}}\right)\right)w^{n}(\underline{x})-\exp\left(-\frac{c}{\lambda_{1}}\right)(\mathcal{M}^{\lambda_{2}}\psi^{\bm{\lambda},n}(\underline{x})-\mathcal{M}^{\lambda_{2}}\psi^{\bm{\lambda},n-1}(\underline{x})).$
Thus,
$\min_{x\in\mathbb{R}}w^{n}(x)\leq\frac{C_{0}}{r+C_{0}}\|\psi^{\bm{\lambda},n}-\psi^{\bm{\lambda},n-1}\|_{L^{\infty}(\mathbb{R})}.$ (A.18)
Combining (A.17) and (A.18) yields
$\|\psi^{\bm{\lambda},n+1}-\psi^{\bm{\lambda},n}\|_{L^{\infty}(\mathbb{R})}\leq\frac{C_{0}}{r+C_{0}}\|\psi^{\bm{\lambda},n}-\psi^{\bm{\lambda},n-1}\|_{L^{\infty}(\mathbb{R})}.$
@@
For simplicity, let $q\coloneqq\frac{C_{0}}{r+C_{0}}$. Fix some $n\in\mathbb{N}^{+}$, for any integer $m>n$, applying the above inequality iteratively, we have
@@ Appendix B Additional Experimental Results
In figs. 3a and 3b, we display the impact of $\sigma$ on the learned value function with respect to two different $\bm{\lambda}$ values. For each $\bm{\lambda}$, the curves retain the same shape and monotone ordering in $\sigma$ (flatter near the origin and higher overall as $\sigma$ increases), reflecting a broader effective continuation region and a reduced marginal value of aggressive impulses. This observation suggests that the impact of $\sigma$ on the randomized value functions enjoys a general consistency across different $\bm{\lambda}$ values. Comparing the two $\bm{\lambda}$ instances, stronger entropy-regularization with $\bm{\lambda}=(1.0,1.0)$ yields a smoother, slightly higher cost profile near the origin, consistent with more conservative interventions; weaker regularization with $\bm{\lambda}=(0.5,0.5)$ permits more aggressive corrections and correspondingly lower minima.
In figs. 3c and 3d, we compare the impact of $\sigma$ on the intervention intensity under the two entropy settings, $\bm{\lambda}=(1.0,1.0)$ (stronger entropy) and $\bm{\lambda}=(0.5,0.5)$ (weaker entropy). Similar the observation for value functions, we also observe the general consistency in the impact of $\sigma$ on the intensity across different values of $\sigma$. For any fixed $\sigma$, the regularization effects under different $\bm{\lambda}$ values are clear. When $\|\bm{\lambda}\|_{\infty}$ is reduced from $1.0$ to $0.5$, we can see that the flat region where the intensity $\pi^{*}$ remains close to 0 visibly enlarges; when the state variable escapes away from the flat region, the increase in the intensity gets more rapid. This suggests that if the regularization effect gets tuned down, the controller will explore less and become more committed to exploiting the continuation/intervention decision.

### v5

@@ A.5 Proof in Section 7
$w^{n}(\overline{x}_{\epsilon})$ $\leq\frac{\exp\left(-\frac{c}{\lambda_{1}}\right)}{r+\exp\left(-\frac{c}{\lambda_{1}}\right)}\|g^{n}\|_{L^{\infty}(\mathbb{R})}+\frac{C_{1}}{r}\epsilon(1+|\overline{x}_{\epsilon}|)$ (A.17)
$\leq\frac{C_{0}}{r+C_{0}}\|g^{n}\|_{L^{\infty}(\mathbb{R})}+C_{2}\epsilon(1+|\overline{x}_{\epsilon}|)$
$=K_{n}+C_{2}\epsilon(1+|\overline{x}_{\epsilon}|),$
for some constant $C_{2}>0$ independent of $n,\epsilon$.
We now show that $\epsilon(1+|\overline{x}_{\epsilon}|)\to 0$ and that $\sup_{\mathbb{R}}w^{n}$ is bounded by $K_{n}$. Denote $m_{\epsilon}:=v^{n,\epsilon}(\overline{x}_{\epsilon})=\max_{x\in\mathbb{R}}v^{n,\epsilon}(x)=w^{n}(\overline{x}_{\epsilon})-\frac{\epsilon}{2}\,\overline{x}_{\epsilon}^{2}.$ Since $\overline{x}_{\epsilon}$ maximizes $v^{n,\epsilon}$, we have $v^{n,\epsilon}(\overline{x}_{\epsilon})\geq v^{n,\epsilon}(0)$, that is,
$w^{n}(\overline{x}_{\epsilon})-\frac{\epsilon}{2}\,\overline{x}_{\epsilon}^{2}\geq w^{n}(0),$
and hence
$\frac{\epsilon}{2}\,\overline{x}_{\epsilon}^{2}\leq w^{n}(\overline{x}_{\epsilon})-w^{n}(0)\leq|w^{n}(\overline{x}_{\epsilon})|+|w^{n}(0)|\leq 2\|w^{n}\|_{L^{\infty}(\mathbb{R})}.$
Thus $\overline{x}_{\epsilon}=O(\epsilon^{-1/2})$, and $\epsilon\,|\overline{x}_{\epsilon}|\to 0$ as $\epsilon\to 0$.
Let $M^{*}:=\sup_{x\in\mathbb{R}}w^{n}(x)$. For each $\epsilon>0$,
$m_{\epsilon}=\max_{x\in\mathbb{R}}v^{n,\epsilon}(x)\leq\sup_{x\in\mathbb{R}}w^{n}(x)=M^{*}.$
Conversely, for any $\eta>0$ there exists $x_{\eta}\in\mathbb{R}$ such that $w^{n}(x_{\eta})\geq M^{*}-\eta$, and hence
$m_{\epsilon}\geq v^{n,\epsilon}(x_{\eta})=w^{n}(x_{\eta})-\frac{\epsilon}{2}x_{\eta}^{2}\geq M^{*}-\eta-\frac{\epsilon}{2}x_{\eta}^{2}.$
Letting $\epsilon\downarrow 0$ for fixed $\eta$ yields $\liminf_{\epsilon\downarrow 0}m_{\epsilon}\geq M^{*}-\eta$, and since $\eta>0$ is arbitrary we conclude that $m_{\epsilon}\to M^{*}$ as $\epsilon\downarrow 0$. Moreover, $w^{n}(\overline{x}_{\epsilon})\geq v^{n,\epsilon}(\overline{x}_{\epsilon})=m_{\epsilon}$, so
$\limsup_{\epsilon\downarrow 0}w^{n}(\overline{x}_{\epsilon})\;\geq\;\lim_{\epsilon\downarrow 0}m_{\epsilon}=M^{*}.$
On the other hand, from (A.17) and $\epsilon(1+|\overline{x}_{\epsilon}|)\to 0$ we obtain
$\limsup_{\epsilon\downarrow 0}w^{n}(\overline{x}_{\epsilon})\leq K_{n}.$
Combining the two displays gives $M^{*}\leq K_{n}$, i.e.
$\sup_{x\in\mathbb{R}}w^{n}(x)\leq K_{n}.$
An entirely analogous argument applied to $w^{n}$ in place of $w^{n}$ gives $\inf_{x\in\mathbb{R}}w^{n}(x)\leq K_{n}$. Therefore
$\|w^{n}\|_{L^{\infty}(\mathbb{R})}=\sup_{x\in\mathbb{R}}|w^{n}(x)|\leq K_{n}=\frac{C_{0}}{r+C_{0}}\,\|\mathcal{M}^{\lambda_{2}}\psi^{\bm{\lambda},n}-\mathcal{M}^{\lambda_{2}}\psi^{\bm{\lambda},n-1}\|_{L^{\infty}(\mathbb{R})},$
which is the desired contraction estimate.
@@
Combining with non-expansiveness of $\mathcal{M}^{\lambda_{2}}$ yields $\|\psi^{\bm{\lambda},n+1}-\psi^{\bm{\lambda},n}\|_{L^{\infty}(\mathbb{R})}\leq\frac{C_{0}}{r+C_{0}}\|\psi^{\bm{\lambda},n}-\psi^{\bm{\lambda},n-1}\|_{L^{\infty}(\mathbb{R})}$. For simplicity, let $q\coloneqq\frac{C_{0}}{r+C_{0}}$. Fix some $n\in\mathbb{N}^{+}$, for any integer $m>n$, applying the above inequality iteratively, we have
@@ Appendix B Additional Experimental Results
In Figures 3a and 3b, we display the impact of $\sigma$ on the learned value function with respect to two different $\bm{\lambda}$ values, i.e., $\bm{\lambda}=(1.0,1.0),(0.5,0.5)$. For each $\bm{\lambda}$, the curves retain the same shape and monotone ordering in $\sigma$ (flatter near the origin and higher overall as $\sigma$ increases), reflecting a broader effective continuation region and a reduced marginal value of aggressive impulses. This observation suggests that the impact of $\sigma$ on the randomized value functions enjoys a general consistency across different $\bm{\lambda}$ values. Comparing the two $\bm{\lambda}$ instances, stronger entropy-regularization with $\bm{\lambda}=(1.0,1.0)$ yields a smoother, slightly higher cost profile near the origin, consistent with more conservative interventions; weaker regularization with $\bm{\lambda}=(0.5,0.5)$ permits more aggressive corrections and correspondingly lower minima.
In Figures 3c and 3d, we compare the impact of $\sigma$ on the intervention intensity under the two entropy settings. Similar to the observation for value functions, we also observe the general consistency in the impact of $\sigma$ on the intensity across different values of $\sigma$. For any fixed $\sigma$, the regularization effects under different $\bm{\lambda}$ values are clear. When regularization is weaker, we can see that the flat region where the intensity $\pi^{*}$ remains close to 0 visibly enlarges, representing a larger waiting region; when the state variable escapes away from the flat region, the increase in the intensity gets more rapid, representing for a more aggressive intervention.